# Real-World Offline Reinforcement Learning from Vision Language Model Feedback

### v1

@@ Abstract
Offline reinforcement learning can enable policy learning from pre-collected, sub-optimal datasets without online interactions. This makes it ideal for real-world robots and safety-critical scenarios, where collecting online data or expert demonstrations is slow, costly, and risky. However, most existing offline RL works assume the dataset is already labeled with the task rewards, a process that often requires significant human effort, especially when ground-truth states are hard to ascertain (e.g., in the real-world). In this paper, we build on prior work, specifically RL-VLM-F, and propose a novel system that automatically generates reward labels for offline datasets using preference feedback from a vision-language model and a text description of the task. Our method then learns a policy using offline RL with the reward-labeled dataset. We demonstrate the system’s applicability to a complex real-world robot-assisted dressing task, where we first learn a reward function using a vision-language model on a sub-optimal offline dataset, and then we use the learned reward to employ Implicit Q learning to develop an effective dressing policy. Our method also performs well in simulation tasks involving the manipulation of rigid and deformable objects, and significantly outperform baselines such as behavior cloning and inverse RL. In summary, we propose a new system that enables automatic reward labeling and policy learning from unlabeled, sub-optimal offline datasets. Videos can be found on our project website [^1] https://sites.google.com/andrew.cmu.edu/offline-rl-vlm-f?usp=sharing.
Figure 1: Top: Our system, Offline RL-VLM-F, combines RL-VLM-F [1] with offline RL for effective policy learning from unlabeled datasets. Given an unlabeled dataset without rewards, it first samples observation pairs, and queries a Vision Language Model for preferences given a text description of the task. Using the preference labels, it then learns a reward model via preference-based reward learning. The learned reward model is utilized to label the entire offline dataset. Finally, it performs offline RL with the labeled dataset to learn a policy that solves the task. Bottom: we follow the same VLM querying process as in RL-VLM-F [1]. It consists of two stages: the first is an analysis stage that asks the VLM to analyze and compare the two images; the second is the labeling stage, where the VLM generates the preference labels based on its own analysis from the first stage and the task description.
@@ I INTRODUCTION
Most prior work assumes that the offline dataset comes with labeled rewards and focus on algorithm design [2, 3, 4, 5]. While reward labeling can be straightforward for simple tasks or those in simulation where the rewards can be generated based on low-level environment states, it becomes significantly more challenging for complex or real-world tasks where ground-truth states are not easily accessible. In such cases, manual reward labeling is often needed, making it a time-consuming bottleneck in applying offline RL algorithms. In this paper, we propose a system that automatically generates reward labels for a given offline dataset, enabling the learning of effective robot control policies from unlabeled, sub-optimal datasets.
@@
In this paper, we adapt RL-VLM-F for the offline RL setting. Specifically, we query a VLM to generate a preference dataset from the given offline dataset. Subsequently, we learn a reward function from the generated preference dataset, and we use the reward function to label the given dataset. The dataset with the labelled rewards can then be utilized with existing frameworks for offline reinforcement learning to learn a control policy. An overview of our system is shown in Figure 1. Furthermore, we demonstrate the effectiveness of the proposed system by applying it to a complex real-world robot-assisted dressing task, learning a point-cloud based reward and policy from a sub-optimal unlabeled dataset.
@@ II Related Work
Behaviour Cloning (BC). Behavior cloning [22] is a classic imitation learning method where an agent learns to perform tasks by mimicking expert demonstrations. The objective in BC is to learn a policy that maps states to actions, often achieved via supervised learning by minimizing the discrepancy between the agent’s and the expert’s actions. BC methods do not require the dataset to contain the reward information, but it usually assumes the demonstrations are optimal in solving the task. Many recent works have shown that Behavior Cloning can be a strong method in learning complex robotic manipulation policies with the right policy representation [23, 24, 25, 26, 27]. Our method uses offline reinforcement learning with the learned reward, and does not require either labelled rewards nor the demonstrations to be optimal.
@@ III Preliminaries
Offline RL and Implicit Q Learning (IQL) [2]. We consider the standard offline RL setting, in which the agent has no access to the online environment interactions but instead learns from a fixed dataset of transitions $D=\{s_{i},a_{i},r_{\psi}(s_{i},a_{i}),s^{\prime}_{i}\}_{i=1}^{N}$ of size $N$. These transitions can be collected with a behaviour policy $\pi_{\beta}$, which might be a mixture of several sub-optimal policies. In contrast to prior work that assumes the reward is given, here we assume the reward is labelled by a learned function $r_{\psi}$, which in our case is learned using preference labels provided by a VLM (more details in Section IV).
We use IQL [2] as the underlying offline RL algorithm for all our experiments, and we briefly review how it works. IQL predicts an upper expectile of the TD-target, which approximates the maximum of $r(s,a)+\gamma Q_{\theta}(s^{\prime},a^{\prime})$ over actions $a^{\prime}$ constrained to the dataset actions. IQL trains a separate value and Q function. The value function $V_{\psi}$ is trained to approximate an expectile purely with respect to the action distribution:
@@ IV-A Reward labeling of the offline dataset
Query VLM: We query the VLM to obtain preferences over the sampled observation pairs. The VLM evaluates the pair based on the provided task description and image observations. The output preference from the VLM is stored as labels. As shown in the bottom part of Fig. 1, we follow the procedure from RL-VLM-F to query the VLM. The querying process involves two stages: analysis and labeling. In the analysis stage, we query the VLM for detailed responses describing and comparing how well each of two images achieves the task goal. In the labeling stage, we use the VLM-generated text responses to extract a preference label between the images. Specifically, we ask the VLM to generate a preference label $y\in\{-1,0,1\}$, where 0 indicates the first image is better, 1 indicates the second image is better, and -1 indicates no discernible differences, based on its own response from the analysis stage. Image pairs labeled as -1 are not used to train the reward model. For both stages, we use a fixed prompt template, and only fill in different text descriptions for different tasks. See Fig. 1 for the prompt templates we used for these two stages.
@@
IV-B Policy Learning from the labelled dataset
In the policy learning phase, we first label the entire offline dataset using the learned reward model. From the labelled dataset, we then learn a policy using Implicit Q Learning.
@@ IV-C Implementation Details
In the real world robot-assisted upper-body dressing task, we don’t have access to the low-level states of the task for two reasons: 1) the cloth is highly deformable and does not have a known compact low-level state; 2) even if we manually define a low-level state as e.g., key points on the cloth, it is challenging to accurately estimate these key points reliably from sensory observations. Therefore, we choose to just represent the cloth, along with the human arm, using point clouds obtained from a depth camera, as was done in prior work [30]. We use this point cloud representation for both learning the reward model and the policy. Note that when querying the VLM, we still provide it with pairs of image observations, because the VLM was trained from images; however; after we store the preference labels from the VLM, we can learn the reward model using the corresponding point cloud observations. In simulation, we learn all the reward models from images, and the policy with low-level ground-truth states provided by the simulator. For the real-world dressing task, both the reward model and the policy are represented using the PointNet++ [36] architecture; for all simulation tasks, the reward model is a ResNet model [37] and the policy is an MLP.
@@
TABLE I: Performance of all compared algorithms in 4 simulation tasks. For each task, we test with 3 datasets of different level of optimalities. For Open Drawer and Soccer, the number represents the success rate; for other tasks the number represents the ground-truth return. The best performing method other than IQL-GT Reward is bolded.
Task Dataset IQL-GT Reward Offline RLVLMF (Ours) IQL-Average Reward Simple Behavioral Cloning Diffusion Policy GAIL
Random 0.99 (0.02) 0.91 (0.08) 0.19 (0.12) 0.04 (0.06) 0.08 (0.06) 0.03(0.05)
Open Drawer Medium 0.95 (0.06) 0.85 (0.13) 0.84 (0.14) 0.88 (0.10) 0.00 (0.01) 0.00(0.00)
Expert 0.99 (0.01) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 0.07(0.09)
Random 0.30 (0.15) 0.23 (0.11) 0.11 (0.09) 0.11 (0.08) 0.13 (0.14) 0.01(0.02)
Soccer Medium 0.19 (0.10) 0.15 (0.10) 0.19 (0.10) 0.20 (0.12) 0.10 (0.12) 0.00(0.01)
Expert 0.42 (0.13) 0.42 (0.12) 0.41 (0.13) 0.41 (0.13) 0.49 (0.17) 0.04(0.06)
Random -93.54 (20.22) -212.87 (36.84) -2537.62 (94.90) -1816.98 (109.75) -2063.65 (89.31) -1683.31(N/A)
Cartpole Medium -98.85 (49.86) -144.56 (75.11) -236.77 (27.70) -165.16 (14.03) -1875.05 (149.51) -431.31(N/A)
Expert -83.93 (14.89) -78.77 (0.83) -81.46 (5.18) -79.19 (0.99) -1423.32 (309.85) -433.64(N/A)
Random 20.93 (0.28) 16.70 (1.12) 8.71 (1.43) 12.73 (1.17) 12.01(3.44) 11.75(0.30)
Straighten Rope Medium 14.34 (1.00) 14.56 (0.96) 14.15 (1.11) 14.17 (0.98) 20.27(1.44) 19.88(1.53)
Expert 20.58 (0.30) 20.54 (0.32) 20.46 (30.83) 20.54 (0.31) 20.27(0.30) 12.33(2.86)
@@ V-A1 Experiment setting
We compare our method with the following baselines: 1. Simple behavior cloning (BC): we perform standard behavior cloning on the dataset to learn the policy. The policy is represented as a simple MLP. 2. GAIL [17]: an inverse RL algorithm that uses adversarial training to learn a cost and policy from the given demonstrations. 3. Diffusion Policy [24]: a state-of-the-art imitation learning algorithm that represents the policy as a denoising diffusion process. 4. IQL with average reward: this baseline labels all transitions in the dataset with the same reward, which is the average of the ground-truth reward of the dataset. It then performs IQL with the labelled dataset. Such a baseline has been proposed and used in prior work [41]. 5. IQL with ground-truth reward: this baseline performs IQL on the dataset, using ground-truth rewards. This baseline is an oracle and provides an upper bound for the performance. Among these baselines, Simple BC, GAIL, and diffusion policy usually assume the dataset to be optimal for learning effective policies.
@@ V-A2 Simulation Results
The results for all simulation experiments can be found in Table I. As shown, for all tasks, when using the random dataset, our method outperforms all compared methods by a large margin (we do not consider the comparison to IQL-GT Reward which uses the ground-truth reward). This demonstrates the effectiveness of the learned reward using VLM preferences, and that our system can produce effective policies from low-quality unlabeled datasets. When using the medium dataset, we find our method to perform the best on Cartpole, and slightly worse than the best performing methods on Open Drawer and Soccer. We also find it interesting that our method often performs better on the Random dataset than on the medium dataset; we hypothesize the reason to be that the medium dataset contains less exploration trajectories and thus less state coverage. As expected, the performance of most methods are best when trained on the expert dataset. We find Diffusion Policy to fail drastically for Cartpole even with expert demonstrations. We speculate the reason is that Cartpole is more dynamic so the velocity information is more important. Finally, we find the learning of GAIL to be unstable and it fails to learn meaningful policies on most tasks, which has been reported in prior work as well [20]. The N/A in GAIL for Cartpole means two seeds run into NaN issues and the learning completely failed.

### v2

@@ Abstract
Offline reinforcement learning can enable policy learning from pre-collected, sub-optimal datasets without online interactions. This makes it ideal for real-world robots and safety-critical scenarios, where collecting online data or expert demonstrations is slow, costly, and risky. However, most existing offline RL works assume the dataset is already labeled with the task rewards, a process that often requires significant human effort, especially when ground-truth states are hard to ascertain (e.g., in the real-world). In this paper, we build on prior work, specifically RL-VLM-F, and propose a novel system that automatically generates reward labels for offline datasets using preference feedback from a vision-language model and a text description of the task. Our method then learns a policy using offline RL with the reward-labeled dataset. We demonstrate the system’s applicability to a complex real-world robot-assisted dressing task, where we first learn a reward function using a vision-language model on a sub-optimal offline dataset, and then we use the learned reward to employ Implicit Q learning to develop an effective dressing policy. Our method also performs well in simulation tasks involving the manipulation of rigid and deformable objects, and significantly outperforms baselines such as behavior cloning and inverse RL. In summary, we propose a new system that enables automatic reward labeling and policy learning from unlabeled, sub-optimal offline datasets. Videos can be found on our project website [^1] https://offline-rlvlmf.github.io/.
Figure 1: Top: Our system, Offline RL-VLM-F, combines RL-VLM-F [1] with offline RL for effective policy learning from unlabeled datasets. Given an unlabeled dataset without rewards, it first samples observation pairs and queries a Vision Language Model for preferences given a text description of the task. Using the preference labels, it then learns a reward model via preference-based reward learning. The learned reward model is utilized to label the entire offline dataset. Finally, it performs offline RL with the labeled dataset to learn a policy that solves the task. Bottom: We follow the same VLM querying process as in RL-VLM-F [1]. It consists of two stages: the first is an analysis stage that asks the VLM to analyze and compare the two images; the second is the labeling stage, where the VLM generates the preference labels based on its own analysis from the first stage and the task description.
@@ I INTRODUCTION
Most prior work assumes that the offline dataset comes with labeled rewards and focuses on algorithm design [2, 3, 4, 5]. While reward labeling can be straightforward for simple tasks or those in simulation where the rewards can be generated based on low-level environment states, it becomes significantly more challenging for complex or real-world tasks where ground-truth states are not easily accessible. In such cases, manual reward labeling is often needed, making it a time-consuming bottleneck when applying offline RL algorithms. In this paper, we propose a system that automatically generates reward labels for a given offline dataset, enabling the learning of effective robot control policies from unlabeled, sub-optimal datasets.
@@
In this paper, we adapt RL-VLM-F for the offline RL setting. Specifically, we query a VLM to generate a preference dataset from the given offline dataset. Subsequently, we learn a reward function from the generated preference dataset, and we use the reward function to label the given dataset. The dataset with the labeled rewards can then be utilized with existing frameworks for offline reinforcement learning to learn a control policy. An overview of our system is shown in Figure 1. Furthermore, we demonstrate the effectiveness of the proposed system by applying it to a complex real-world robot-assisted dressing task, learning a point-cloud based reward and policy from a sub-optimal unlabeled dataset.
@@ II Related Work
Behavior Cloning (BC). Behavior cloning [22] is a classic imitation learning method where an agent learns to perform tasks by mimicking expert demonstrations. The objective in BC is to learn a policy that maps states to actions, often achieved via supervised learning by minimizing the discrepancy between the agent’s and the expert’s actions. BC methods do not require the dataset to contain the reward information, but they usually assume the demonstrations are optimal in solving the task. Many recent works have shown that Behavior Cloning can be a strong method in learning complex robotic manipulation policies with the right policy representation [23, 24, 25, 26, 27]. Our method uses offline reinforcement learning with the learned reward and does not require either labeled rewards or demonstrations to be optimal.
@@ III Preliminaries
Offline RL and Implicit Q Learning (IQL) [2]. We consider the standard offline RL setting, in which the agent has no access to the online environment interactions but instead learns from a fixed dataset of transitions $D=\{s_{i},a_{i},r_{\psi}(s_{i},a_{i}),s^{\prime}_{i}\}_{i=1}^{N}$ of size $N$. These transitions can be collected with a behavior policy $\pi_{\beta}$, which might be a mixture of several sub-optimal policies. In contrast to prior work that assumes the reward is given, here we assume the reward is labeled by a learned function $r_{\psi}$, which in our case is learned using preference labels provided by a VLM (more details in Section IV).
For all our experiments, we adopt Implicit Q-Learning (IQL)[2] as the base offline RL algorithm. IQL predicts an upper expectile of the TD-target, which approximates the maximum of $r(s,a)+\gamma Q_{\theta}(s^{\prime},a^{\prime})$ over actions $a^{\prime}$ constrained to the dataset actions. IQL trains a separate value and Q function. The value function $V_{\psi}$ is trained to approximate an expectile purely with respect to the action distribution:
@@ IV-A Reward labeling of the offline dataset
Query VLM: We query the VLM to obtain preferences over the sampled observation pairs. The VLM evaluates the pair based on the provided task description and image observations. The output preference from the VLM is stored as labels. As shown in the bottom part of Fig. 1, we follow the procedure from RL-VLM-F to query the VLM. The querying process involves two stages: analysis and labeling. In the analysis stage, we query the VLM for detailed responses describing and comparing how well each of two images achieves the task goal. In the labeling stage, we use the VLM-generated text responses to extract a preference label between the images. Specifically, we ask the VLM to generate a preference label $y\in\{-1,0,1\}$, where 0 indicates the first image is better, 1 indicates the second image is better, and -1 indicates no discernible differences, based on its own response from the analysis stage. Image pairs labeled as -1 are not used to train the reward model. For both stages, we use a fixed prompt template and only fill in different text descriptions for different tasks. See Fig. 1 for the prompt templates we used for these two stages.
@@
IV-B Policy Learning from the labeled dataset
In the policy learning phase, we first label the entire offline dataset using the learned reward model. From the labeled dataset, we then learn a policy using Implicit Q Learning.
@@ IV-C Implementation Details
In the real world robot-assisted upper-body dressing task, we do not have access to the low-level states of the task for two reasons: 1) the cloth is highly deformable and does not have a known compact low-level state; 2) even if we manually define a low-level state as e.g., key points on the cloth, it is challenging to accurately estimate these key points reliably from sensory observations. Therefore, we choose to just represent the cloth, along with the human arm, using point clouds obtained from a depth camera, as was done in prior work [30]. We use this point cloud representation for both learning the reward model and the policy. Note that when querying the VLM, we still provide it with pairs of image observations, because the VLM was trained from images; however; after we store the preference labels from the VLM, we can learn the reward model using the corresponding point cloud observations. In simulation, we learn all the reward models from images, and the policy with low-level ground-truth states provided by the simulator. For the real-world dressing task, both the reward model and the policy are represented using the PointNet++ [36] architecture; for all simulation tasks, the reward model is a ResNet model [37] and the policy is an MLP. All our experiments were conducted on 1 Nvidia RTX 3090/4090.
@@
TABLE I: Performance of all compared algorithms in 4 simulation tasks. For each task, we test with 3 datasets of different levels of optimalities. For Open Drawer and Soccer, the number represents the success rate; for other tasks the number represents the ground-truth return. The best performing method other than IQL-GT Reward is bolded.
Task Dataset
IQL
GT reward
Offline
RL-VLM-F (Ours)
IQL
Avg. Reward
BC Diffusion Policy GAIL CLIP Reward
Random 0.99 (0.02) 0.91 (0.08) 0.19 (0.12) 0.04 (0.06) 0.08 (0.06) 0.03 (0.05) 0.00 (0.00)
Open
Drawer
Medium 0.95 (0.06) 0.85 (0.13) 0.84 (0.14) 0.88 (0.10) 0.00 (0.01) 0.00 (0.00) 0.24 (0.11)
Expert 0.99 (0.01) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 0.07 (0.09) 0.99 (0.00)
Random 0.30 (0.15) 0.23 (0.11) 0.11 (0.09) 0.11 (0.08) 0.13 (0.14) 0.01 (0.02) 0.15 (0.01)
Soccer Medium 0.19 (0.10) 0.15 (0.10) 0.19 (0.10) 0.20 (0.12) 0.10 (0.12) 0.00 (0.01) 0.16 (0.02)
Expert 0.42 (0.13) 0.42 (0.12) 0.41 (0.13) 0.41 (0.13) 0.49 (0.17) 0.04 (0.06) 0.28 (0.06)
Random -93.54 (20.22) -212.87 (36.84) -2537.62 (94.90) -1816.98 (109.75) -2063.65 (89.31) -1683.31 (N/A) -2672.38 (3.07)
Cartpole Medium -98.85 (49.86) -144.56 (75.11) -236.77 (27.70) -165.16 (14.03) -1875.05 (149.51) -431.31 (N/A) -270.32 (9.28)
Expert -83.93 (14.89) -78.77 (0.83) -81.46 (5.18) -79.19 (0.99) -1423.32 (309.85) -433.64 (N/A) -79.69 (0.23)
Random 20.93 (0.28) 16.70 (1.12) 8.71 (1.43) 12.73 (1.17) 12.01 (3.44) 11.75 (0.30) 13.66 (0.75)
Straighten
Rope
Medium 14.34 (1.00) 14.56 (0.96) 14.15 (1.11) 14.17 (0.98) 20.27 (1.44) 19.88 (1.53) 14.17 (0.21)
Expert 20.58 (0.30) 20.54 (0.32) 20.46 (30.83) 20.54 (0.31) 20.27 (0.30) 12.33 (2.86) 20.49 (0.08)
@@ V-A1 Experiment setting
We compare our method with the following baselines: 1. Simple behavior cloning (BC): we perform standard behavior cloning on the dataset to learn the policy. The policy is represented as a simple MLP. 2. GAIL [17]: an inverse RL algorithm that uses adversarial training to learn a cost and policy from the given demonstrations. 3. Diffusion Policy [24]: a state-of-the-art imitation learning algorithm that represents the policy as a denoising diffusion process. 4. IQL with average reward: this baseline labels all transitions in the dataset with the same reward, which is the average of the ground-truth reward of the dataset. It then performs IQL with the labelled dataset. Such a baseline has been proposed and used in prior work [41]. 5. CLIP reward [12], which computes the reward as the cosine similarity score between the CLIP [42] embedding of the task description and the image observation. 6. IQL with ground-truth reward: this baseline performs IQL on the dataset, using ground-truth rewards. This baseline is an oracle and provides an upper bound for the performance. Among these baselines, Simple BC, GAIL, and diffusion policy usually assume the dataset to be optimal for learning effective policies.
@@ V-A2 Simulation Results
The results for all simulation experiments can be found in Table I. As shown, for all tasks, when using the random dataset, our method outperforms all compared methods by a large margin (we do not consider the comparison to IQL-GT Reward, which uses the ground-truth reward). This demonstrates the effectiveness of the learned reward using VLM preferences and that our system can produce effective policies from low-quality unlabeled datasets. When using the medium dataset, we find our method performs the best on Cartpole, and slightly worse than the best-performing methods on Open Drawer and Soccer. We also find it interesting that our method often performs better on the Random dataset than on the medium dataset; we hypothesize the reason to be that the medium dataset contains less exploration trajectories and thus less state coverage. As expected, the performance of most methods is best when trained on the expert dataset. We find Diffusion Policy to fail drastically for Cartpole even with expert demonstrations. We speculate the reason is that Cartpole is more dynamic, so the velocity information is more important. Finally, we find the learning of GAIL to be unstable, and it fails to learn meaningful policies on most tasks, which has been reported in prior work as well [20]. The N/A in GAIL for Cartpole means two seeds run into NaN issues and the learning completely failed.
V-A3 Ablation Study
VLMs have errors when giving the preference label. We perform ablation studies to examine how robust our system is to errors in the preference labels. Specifically, we perform IQL with preference label obtained using the ground-truth reward, but with probability $p$ we flip the label of the pair. We test with $p=0,0.25,0.5,0.75$. In addition, prior work [1] has found that the error rate of the VLM increases as the underlying states of the two images are closer to each other. Therefore, we also test the case where the probability $p$ of flipping the label is inversely proportional to how close the two states are in the pair, termed ”proximity flipping”. The results are summarized in Table II. As shown, our system is robust to $25\%$ of the error in the preference labels. As the error rate $p$ increases, the performance decreases. As expected, given an error rate $p$, the performance is worse on dataset with lower quality (e.g., the Random dataset). The proximity flipping results are slightly worse than that of having an error rate of 50% but better than $75\%$.
TABLE II: We study the performance of our method across a varying quality of preference datasets. Columns 0.25, 0.5, 0.75 represent the probability of the label being incorrect. The proximity-based flip is when the probability of the label being incorrect is proportional to how close the states are.
Task Dataset IQL-GT Reward 0.25 0.5 0.75 Proximity based Flip
Random 0.99 (0.02) 0.99 (0.02) 0.51 (0.10) 0.00 (0.00) 0.00 (0.01)
Open Drawer Medium 0.95 (0.06) 0.52 (0.17) 0.16 (0.13) 0.01 (0.00) 0.02 (0.01)
Expert 0.99 (0.01) 0.99 (0.00) 0.99 (0.00) 0.92 (0.04) 0.96 (0.05)
Random 0.30 (0.15) 0.15 (0.01) 0.15 (0.01) 0.15 (0.01) 0.19 (0.04)
Soccer Medium 0.19 (0.10) 0.20 (0.04) 0.17 (0.03) 0.15 (0.03) 0.21 (0.03)
Expert 0.42 (0.13) 0.26 (0.06) 0.30 (0.08) 0.28 (0.01) 0.28 (0.05)
Random -93.54 (20.22) -89.72 (0.77) -2677.65 (1.86) -2381.72 (24.71) -214.92 (8.19)
CartPole Medium -98.85 (49.86) -99.16 (5.31) -259.92 (6.35) -773.81 (135.11) -650.91 (147.03)
Expert -83.93 (14.89) -78.61 (0.10) -79.72 (0.24) -117.26 (28.18) -85.69 (11.48)
Random 20.93 (0.28) 18.07 (0.39) 14.63 (1.26) 9.19 (0.45) 17.61 (0.31)
Straighten Rope Medium 14.34 (1.00) 14.33 (0.28) 14.37 (0.16) 14.27 (0.28) 14.05 (0.16)
Expert 20.58 (0.30) 20.54 (0.04) 20.41 (0.04) 20.50 (0.07) 20.51 (0.03)